# Targeted Unlearning Using Perturbed Sign Gradient Methods With Applications On Medical Images

### v1

@@ 2.2 Unlearning as Bilevel Optimization
[…]
@@ 2.4 Outer Optimization
[…]
@@ 2.6 Model Training
All training is performed using cross-entropy loss and SGD with momentum 0.9, a learning rate of $1\times 10^{-3}$, and $L_{2}$ regularization with coefficient $1\times 10^{-4}$. Data augmentation includes random horizontal flips, random rotations up to $15^{\circ}$, color jittering, and random cropping. An 80/20 train–test split is used across all datasets. Training was terminated once the validation accuracy had converged (difference between training accuracy of previous epoch and current epoch is $<$ 1, and the model with the highest validation accuracy within (latest) 5 epochs before convergence was chosen for testing purposes.
@@ 2.7 Evaluation Metrics
We evaluate unlearning performance using four primary criteria. Forget Set Accuracy (F-Acc) measures the accuracy on a held-out portion of the forget set $F$; lower values indicate more effective forgetting. Remain Set Accuracy (R-Acc) captures the accuracy on a held-out portion of the remain set $R$; higher values reflect better retention of model utility. To assess the balance between these objectives, we compute the F/R Accuracy Ratio, defined as the ratio of forget to remain accuracy; lower values suggest better selective forgetting without compromising performance on $R$. Finally, we measure vulnerability to Membership Inference Attacks (MIA) [40] following the protocol of [6]. A logistic regression classifier is trained to distinguish between samples in $F$ and a held-out test set drawn from the same class. The attack model takes as input the clipped loss $\mathcal{L}(f_{w}(x),y)$, truncated to the range $[-400,400]$. We use 5-fold cross-validation and report average classification accuracy, where a perfect defense achieves 50%. Together, these metrics quantify whether forgetting was successful (C1), model utility was preserved (C2), and whether post-unlearning privacy risk was mitigated.
@@ 3.2 Clinical Unlearning
Table 1: Performance of unlearning algorithms on clinical datasets. “Camera” (CFP-Clinic) refers to forgetting images captured with a specific imaging device, while $VPF>12$” (Oculoplastic) denotes unlearning based on a clinical threshold for vertical palpebral fissure. R Acc = Remain set accuracy (higher is better), F Acc = Forget set accuracy (lower is better), F/R = Forget-to-Remain ratio (lower is better), and MIA = Membership Inference Attack accuracy (mean ± std). Bolded values indicate best performance column-wise.
@@ F Acc R Acc
$\boldsymbol{\boldsymbol{-}}$${\boldsymbol{-}}$${\boldsymbol{-}}$ 0.094 0.81
$\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$${\boldsymbol{-}}$ 0.48 0.85
$\boldsymbol{\boldsymbol{-}}$ $\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$ 0.062 0.17
$+\boldsymbol{\boldsymbol{+}}$$\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$ 0.062 0.17
$\boldsymbol{\boldsymbol{+}}$(80) $\boldsymbol{\boldsymbol{+}}$(80) $\boldsymbol{\boldsymbol{-}}$ 0.062 0.17
$\boldsymbol{\boldsymbol{+}}$(95) $\boldsymbol{\boldsymbol{+}}$(95) $\boldsymbol{\boldsymbol{-}}$ 0.062 0.17
$+\boldsymbol{\boldsymbol{+}}$$\boldsymbol{\boldsymbol{+}}$ $\boldsymbol{\boldsymbol{+}}$ 0.51 0.89
@@ Appendix A Appendix
$-\mathcal{L}(\delta)=-\frac{1}{2u}\|\delta\|^{2}_{2}+\mathcal{L}(f_{k}(\delta+x_{i},w_{0})).$
@@
First, we will use $l$ to denote $-L$. So the inner optimization becomes the following minimization problem,
@@
Summing up the inequalities in (16) from $t=0$ to $t=T-1$ we get,
@@
where we used the fact $\delta_{0}$ is initialized at 00, and that $\left(\mathbb{E}[l(\delta^{T})]-\inf l\right)\geq 0$. By choosing $\gamma$ to be the coordinatewise minimum value of gradient magnitude $|\nabla l(\delta^{t})|$, we can assure that $\left|\nabla l(\delta^{t})\right|^{\top}\operatorname{erf}\left(\frac{\left|\nabla l(\delta^{t})\right|}{\gamma\sqrt{2}}\right)=c\|\nabla l\|_{1},$ and the claim follows by following similar techniques as in [47] Theorem 5.12, since the step size/learning rate $\epsilon_{t}=1/t$.
@@ A.5.1 Datasets
Table 3: Open source data details. For DR, 260 images from the Indian Diabetic Retinopathy Image Dataset [31] with severity of $>2$ were chosen, and 240 images having a diagnosis of only DR from the Retinal Fundus Multi-Disease Image dataset (RFMID) [33] were selected. For Glaucoma, 168 images from the Online Retinal Fundus Image Database for Glaucoma [32] and 296 images from the G1020 [34] dataset were selected. ‘Other’ images were selected by randomly sampling 500 images from RFMID that did not have DR or Glaucoma as a label. These images have diagnoses of AMD, macular hole, and more [34].
@@ A.5.2 Additional Unlearning
Table 4: Evaluation of the influence of logits and incorporation of the loss from R in unlearning. A $\phi$ value of $41e-$ 4 was used. A number after the plus sign in parentheses denotes the epoch at which the logits were incorporated during outer optimization.
@@ F Acc R Acc
$\boldsymbol{\boldsymbol{-}}$${\boldsymbol{-}}$${\boldsymbol{-}}$ 0.089 0.12 0.094 0.84 0.81 0.81
$\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$${\boldsymbol{-}}$ 0.49 0.52 0.48 0.87 0.85 0.86
$\boldsymbol{\boldsymbol{-}}$ $\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$+\boldsymbol{\boldsymbol{+}}$$\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$\boldsymbol{\boldsymbol{+}}$(80) $\boldsymbol{\boldsymbol{+}}$(80) $\boldsymbol{\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$\boldsymbol{\boldsymbol{+}}$(90) $\boldsymbol{\boldsymbol{+}}$(90) $\boldsymbol{\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$+\boldsymbol{\boldsymbol{+}}$$\boldsymbol{\boldsymbol{+}}$ $\boldsymbol{\boldsymbol{+}}$ 0.44 0.49 0.49 0.88 0.87 0.88
Table 5: Evaluation of the influence of logits and incorporation of the loss from R in unlearning. A $\phi$ value of $21e-$ 2 was used. A number after the plus sign in parentheses denotes the epoch at which the logits were incorporated during outer optimization.
@@
$\boldsymbol{\boldsymbol{-}}$${\boldsymbol{-}}$${\boldsymbol{-}}$ 0.089 0.12 0.094 0.84 0.81 0.81
$\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$${\boldsymbol{-}}$ 0.49 0.52 0.48 0.87 0.85 0.85
$\boldsymbol{\boldsymbol{-}}$ $\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$+\boldsymbol{\boldsymbol{+}}$$\boldsymbol{\boldsymbol{+}}$${\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$\boldsymbol{\boldsymbol{+}}$(80) $\boldsymbol{\boldsymbol{+}}$(80) $\boldsymbol{\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$\boldsymbol{\boldsymbol{+}}$(90) $\boldsymbol{\boldsymbol{+}}$(90) $\boldsymbol{\boldsymbol{-}}$ 0.001 0.047 0.062 0.18 0.17 0.17
$+\boldsymbol{\boldsymbol{+}}$$\boldsymbol{\boldsymbol{+}}$ $\boldsymbol{\boldsymbol{+}}$ 0.52 0.55 0.51 0.89 0.88 0.89
@@ A.5.4 Hyperparameter Tuning
Figure 10: Unlearning results using various combinations of $\gamma$ and $\lambda$ on non-medical imaging datasets. Column 1 denotes the accuracy of all combinations on R, Column 2 Denotes the accuracy of all combinations on F, and Column 3 denotes the Euclidean distance of $[F_{acc},R_{acc}]$ obtained using all combinations of $\gamma$ and $\lambda$ from $[F_{acc},R_{acc}]$ obtained using the retrained model.
Figure 11: Results of unlearning an entire class using various combinations of $\gamma$ and $\lambda$ on medical imaging datasets. Column 1 denotes the accuracy of all combinations on R, Column 2 Denotes the accuracy of all combinations on F, and Column 3 denotes the Euclidean distance of $[F_{acc},R_{acc}]$ obtained using all combinations of $\gamma$ and $\lambda$ from $[F_{acc},R_{acc}]$ obtained using the retrained model. OP denotes oculoplastic dataset, and US denotes ultrasound dataset. Directionality of arrows indicates whether high or low values are the desired output.
Figure 12: Results of unlearning subsets of images based on clinical parameters using various combinations of $\gamma$ and $\lambda$ on medical imaging datasets. Column 1 denotes the accuracy of all combinations on R, Column 2 Denotes the accuracy of all combinations on F, and Column 3 denotes the Euclidean distance of $[F_{acc},R_{acc}]$ obtained using all combinations of $\gamma$ and $\lambda$ from $[F_{acc},R_{acc}]$ obtained using the retrained model. A-C represents the results of unlearning images acquired with a specific imaging device, and D-F are the results of unlearning images of eyes that had a vertical palpebral fissure ¿ 11mm. Directionality of arrows indicates whether high or low values are the desired output.

### v2

@@ 2.2 Unlearning as Bilevel Optimization
$\text{s.t.}\quad x_{i}^{b}-x_{i}\in\arg\min_{\delta:f_{w_{0}}(x_{i}+\delta)^{\top}y_{i}\leq\kappa}\ell(\delta)\quad\forall i\in F,$ (1)
@@ 2.4 Outer Optimization
Second, we allow the outer loss to be computed using dense logits instead of hard labels. To interpolate between sparse (argmax) and dense supervision, we define a simple top-$k$ logit preprocessing step. In all experiments, $k$ was set to 3. In detail, let $f_{w}(x_{i})\in\mathbb{R}^{K}$ denote the logits for sample $x_{i}$, and let $\text{Top-}k(f_{w}(x_{i}))$ denote the indices of the largest $k$ logits. We define a sparse logit vector $\tilde{f}_{w}(x_{i})$ as:
@@ 2.6 Model Training
All training is performed using cross-entropy loss and SGD with momentum 0.9, a learning rate of $1\times 10^{-3}$, and $L_{2}$ regularization with coefficient $1\times 10^{-4}$. Data augmentation includes random horizontal flips, random rotations up to $15^{\circ}$, color jittering, and random cropping. An 80/20 train–test split is used across all datasets. Training was terminated once the validation accuracy had converged (difference between training accuracy of previous epoch and current epoch is << 1, and the model with the highest validation accuracy within (latest) 5 epochs before convergence was chosen for testing purposes.
@@ 2.7 Evaluation Metrics
We evaluate unlearning performance using four primary criteria. Forget Set Accuracy (F-Acc) measures the accuracy on a held-out portion of the forget set $F$; lower values indicate more effective forgetting. Remain Set Accuracy (R-Acc) captures the accuracy on a held-out portion of the remain set $R$; higher values reflect better retention of model utility. To assess the balance between these objectives, we compute the F/R Accuracy Ratio, defined as the ratio of forget to remain accuracy; lower values suggest better selective forgetting without compromising performance on $R$. Finally, we measure vulnerability to Membership Inference Attacks (MIA) [40] following the protocol of [6]. A logistic regression classifier is trained to distinguish between samples in $F$ and a held-out test set drawn from the same class. The attack model takes as input the clipped loss $\mathcal{L}(f_{w}(x),y)$, truncated to the range [400,400][-400,400]. We use 5-fold cross-validation and report average classification accuracy, where a perfect defense achieves 50%. Together, these metrics quantify whether forgetting was successful (C1), model utility was preserved (C2), and whether post-unlearning privacy risk was mitigated.
@@ 3.2 Clinical Unlearning
Table 1: Performance of unlearning algorithms on clinical datasets. “Camera” (CFP-Clinic) refers to forgetting images captured with a specific imaging device, while VPF>12VPF>12” (Oculoplastic) denotes unlearning based on a clinical threshold for vertical palpebral fissure. R Acc = Remain set accuracy (higher is better), F Acc = Forget set accuracy (lower is better), F/R = Forget-to-Remain ratio (lower is better), and MIA = Membership Inference Attack accuracy (mean ± std). Bolded values indicate best performance column-wise.
@@ F Acc R Acc
[…]
@@ Appendix A Appendix
$\mathcal{L}(\delta)=-\frac{1}{2u}\|\delta\|^{2}_{2}+\mathcal{L}(f_{k}(\delta+x_{i},w_{0})).$
@@
First, we will use $l$ to denote L-L. So the inner optimization becomes the following minimization problem,
@@
Summing up the inequalities in (16) from $t=0$ to t=T1t=T-1 we get,
@@
where we used the fact $\delta_{0}$ is initialized at 0, and that $\left(\mathbb{E}[l(\delta^{T})]-\inf l\right)\geq 0$. By choosing $\gamma$ to be the coordinatewise minimum value of gradient magnitude $|\nabla l(\delta^{t})|$, we can assure that $\left|\nabla l(\delta^{t})\right|^{\top}\operatorname{erf}\left(\frac{\left|\nabla l(\delta^{t})\right|}{\gamma\sqrt{2}}\right)=c\|\nabla l\|_{1},$ and the claim follows by following similar techniques as in [47] Theorem 5.12, since the step size/learning rate $\epsilon_{t}=1/t$.
@@ A.5.1 Datasets
Table 3: Open source data details. For DR, 260 images from the Indian Diabetic Retinopathy Image Dataset [31] with severity of >2>2 were chosen, and 240 images having a diagnosis of only DR from the Retinal Fundus Multi-Disease Image dataset (RFMID) [33] were selected. For Glaucoma, 168 images from the Online Retinal Fundus Image Database for Glaucoma [32] and 296 images from the G1020 [34] dataset were selected. ‘Other’ images were selected by randomly sampling 500 images from RFMID that did not have DR or Glaucoma as a label. These images have diagnoses of AMD, macular hole, and more [34].
@@ A.5.2 Additional Unlearning
Table 4: Evaluation of the influence of logits and incorporation of the loss from R in unlearning. A $\phi$ value of 1e41e-4 was used. A number after the plus sign in parentheses denotes the epoch at which the logits were incorporated during outer optimization.
@@ F Acc R Acc
[…]
Table 5: Evaluation of the influence of logits and incorporation of the loss from R in unlearning. A $\phi$ value of 1e21e-2 was used. A number after the plus sign in parentheses denotes the epoch at which the logits were incorporated during outer optimization.
@@
[…]
@@ A.5.4 Hyperparameter Tuning
Figure 10: Unlearning results using various combinations of $\gamma$ and $\lambda$ on non-medical imaging datasets. Column 1 denotes the accuracy of all combinations on R, Column 2 Denotes the accuracy of all combinations on F, and Column 3 denotes the Euclidean distance of $F_{acc},R_{acc}]$ obtained using all combinations of $\gamma$ and $\lambda$ from $F_{acc},R_{acc}]$ obtained using the retrained model.
Figure 11: Results of unlearning an entire class using various combinations of $\gamma$ and $\lambda$ on medical imaging datasets. Column 1 denotes the accuracy of all combinations on R, Column 2 Denotes the accuracy of all combinations on F, and Column 3 denotes the Euclidean distance of $F_{acc},R_{acc}]$ obtained using all combinations of $\gamma$ and $\lambda$ from $F_{acc},R_{acc}]$ obtained using the retrained model. OP denotes oculoplastic dataset, and US denotes ultrasound dataset. Directionality of arrows indicates whether high or low values are the desired output.
Figure 12: Results of unlearning subsets of images based on clinical parameters using various combinations of $\gamma$ and $\lambda$ on medical imaging datasets. Column 1 denotes the accuracy of all combinations on R, Column 2 Denotes the accuracy of all combinations on F, and Column 3 denotes the Euclidean distance of $F_{acc},R_{acc}]$ obtained using all combinations of $\gamma$ and $\lambda$ from $F_{acc},R_{acc}]$ obtained using the retrained model. A-C represents the results of unlearning images acquired with a specific imaging device, and D-F are the results of unlearning images of eyes that had a vertical palpebral fissure ¿ 11mm. Directionality of arrows indicates whether high or low values are the desired output.